06/11/07 – TE received the final pay application along with the allowance clarification and close-out paperwork. All materials have been sent to HCFD.

04/27/07 – Mark Donnelly has directed TE to let QCE know that they should send their final pay application so that this project can be closed out through TE. Hamilton County Facilities with deal directly with American Sound to complete the remaining PA system work. TTC reviewing EDi Closeout for submissions to County.

04/13/07 – American Sound will directly quote project to revise PA system wiring to Mark Donnelly of Hamilton County Facilities Department. QCE will be closed out of their involvement with this portion of work.

04/05/07 – QCE is working

03/30/07 – Project reported to have received occupancy from City. Bulletin issued for PA system. Final paperwork from QCE after change order is needed.

03/23/07 – Space is occupied and there a couple of items left to finish and closeout. The PA system has to modified for client satisfaction. TE and MD are working with QCE to resolve the issue. TE and MD agreed on which QCE change order will be paid from the allowance (via email) and what will be a change order to the project (extra conduits).

03/20/07 – Punchlist are complete. Everything is working except the paging system.

03/15/07 – We received nice kudos from Frtizz on this job. Closeout meeting went well. County agreed to insulated the steam piping behind the radiator. TTC and Honeywell agreed that some control tweaking still remains to be done before project is completed. TTC received O&M’s from EDI last Monday and must turn them over after review to MD. Once payapps are received and O&M’s approved this project can be closed out for EDI and QC M.

02/23/07 – TEI distributed the mechanical punchlist for this project to EDI and partners. EDI forwarded to Honeywell and requested everything be completed by 02/28/07 for project completion. PDF

02/15/07 – EDI provided an update on the project as they are trying to do their balance report but the unit has no power yet and the controls are not hooked up. QCE was contacted about having power to the unit and was suppose to do that work yesterday as was Honeywell who did not make it down until today. QCE reported that they did have power to the unit but the fusible disconnect on the 5th needed to be turned on except that no one from QCE was there to turn that on. The balancer is on site today waiting for the system to be turned on to get the readings they need to startup the unit and get the balance report complete. EDI sent in a list of change order and allowance request that TTC processed and forwarded to Mark Donnelly for discussion. This list included the condensate pump, relocation of CHWS/R, steam lines, and fire protection supports. MD and TTC discussed by phone and the County is willing to pay at this point for everything but the relocation of the CHWS/R lines.

01/22/07 – Everyone met onsite today to discuss the MER. Space was allocated for the Controls, Data and Mechanical Systems. Electrical coordinated the locations of the additional power outlet for the mechanical sump pump and the receptacle for the controls. Owner stated that fire support piping will be hung from deck and not form interstitial ceiling in Room 468A. TTC showed EDI this space and EDI will adjust the hangers. Owner will not accept ceiling hung from ductwork. Radiators are complete. In reading room TE asked for Architect to check height of new cover to handle the concrete brick under the legs. Steam system is up and running. 3 of the 4 steam control valves are replaced. One could not be removed because of pipe rupture concerns so TE decided to leave as is. Honeywell submitted a 3-way valve for control on the job but that is not required and it has been plugged to a 2-way valve per TTC’s instruction. The new drain and chilled water piping will be installed above the unit at about seven foot AFF and drop down to the unit next to the unit leaving the unit open for access. We will trade the access door work by the GC for a square opening and two backboards for controls and data in the room. TE coordinated these locations and sizes with CRA and the contractors.

01/17/07 – TEI requested that EDI relocate the CHWS/R away from the disconnects and route over top the unit. QCE and CRA decided that the soffit can be lowered 4″. We will issue bulletins for the additional work and get prices for the County. Honeywell reported that their control panel will only require 8A of service which means it can be fed from our AHU unit or electric heater with a separate transformer and given the low IT requirement, it may share the same circuit as it. TEI called for a coordination mtg on MER on 01/22/07 at 1:00 p.m. to coordinated all these issues.

01/16/07 – TEI is meeting QCE to discuss additional conduits under the soffitt in the CH hallway. TEI invited CRA and HamCo to attend on Tuesday at 4pm.

01/10/07 – Sprinkler contractor was on site working on the sprinkler system on 01/09. HVAC work is progressing. TTC instructed EDI to use the condensate pump provided with the unit, that the integral check valve meets our requirements and to use the 1.25″ piping as called for on the dwgs to drain system. EDI asked if the 3-way valve provided by honeywell could be plugged to be a 2-Way valve and TTC stated that this was okay. EDI reported the the GC would be reworking the ceiling grid to match the dwgs. It is currently running in the wrong direction. The sprinkler contractor have the sprinkler mains ready for final arm overs down to the new ceiling and will come back after ceiling is reworked. EDI inquired today about placing the radiator on concrete blocks, I am not sure this is an idea I am going to approve.

01/02/07 – The AHU was delivered into the MER on 12/29/06. EDI will be working nights to hook up the HVAC to the new unit runs. Honeywell will be in to hook up and download the controls soon but an exact timeline has not been set. TTC asked the County to allow us to use $600 for a steam condensate fix for the old air handlers located in Room 468. MD has given TEI approval to do this work under the $2500 allowance for this project.

12/21/06 – Const Mtg # 2 was held today. EDI has VAV boxes installed, all ductwork hung except for MER. EDI stated FD will be installed when unit is installed. Unit is scheduled to be arrive on December 27. QCE needs to communicate better with field visits

12/14/06 – TEI was onsite today to review mechanical and electrical progress. On the Mechanical side EDI has all the main ductwork installed up to the mechanical room. The VAV boxes are installed and the sprinkler work is installed except for the arm overs that will eventually be connected after the ceiling is installed. The exhaust fan is installed but the new louver is not installed yet but will be next on the agenda. We are waiting for the unit to arrive before more work is completed. We discussed the new radiator location in the small reading room and TEI instructed EDI to try to fit the existing radiator in between the piping and use piping crosses to connect before moving the radiator up or more into the room. They will try this and advise. I do not believe that any controls programming has been installed yet. TEI is meeting with QCE and Mitch on site at 4pm to go over data cabling, power routes and the shutdown of power today for our tie-in for the new AHU unit. Mike Bartlett is keeping an electrician late today to assist in case QCE runs into trouble.

12/04/06 – Work is progressing. EDI has submitted a full schedule which I believe will be incorporated by the GC into the master schedule. EDI informs me that the VAV boxes have been delivered to the site. The unit is still scheduled for 12/28. The project schedule is still valid.

11/21/06 – We held a construction and schedule meeting with the contractors. EDI has began work in the space on the mechanical side. Demo has been completed. The old AHU has been removed along with its ductwork. Holes have been cut into the walls for the new ductwork paths. Equipment is ordered and scheduled to be here mid to late December. Electrically, 3/4″ conduits have been snake across the ceiling including the voice and data for HamCo communications. One conduit in the mechanical opening for the new supply ductwork has to be moved and will be by QCE. QCE cut a hole in the soffit to facilitate new conduit paths. The GC will have to patch this hole at the end of the job. It was discussed and no access panel is needed at this location. TEI stated that VAV box has SCR control on electric heat and this may be done through a PID loop by Honeywell only if they can provide enough stages of heating to keep the space a constant temperature. Fire dampers were discussed with the code official and they are not required but an engineering dwg would be required to change them so it isn’t worth the savings. The FD will be installed.

10/27/06 – Construction has started. Light fixture submittals have been approved. TEI approved final AHU submittals with EDI and Trane. The units size will be decreased by 1 size allowing it to fit into the space but this increased the face velocity through the coil to slight above 500 fpm (a normal max for airflow). Trane assured us that there would be no moisture drift through this unit. All equipment is on order.

10/01/06 – Pre Construction meeting today. EDI turned over submittals. ThermalTech reviewed and sent back comments. Trane AHU is revise and resubmit. EF, Louver, Air Devices are all furnish as corrected. PL is talking to QCE about down sizing to two 3/4″ conduits across the hallway. Mitch requested a conduit for the IT but we also discussed leaving room on the wall in the MER for a patch panel and have one cable go back across the hallway for easier access. The county has already remediated the space so it is ready for work. TTC instructed EDI to have sprinkler piping in MER room ran close to the outside wall to stay away from new electrical disconnect switches in the space. EDI is supposed to find out who the FPC is on this job and let TTC know.

09/01/06 – TTC contacted the City Plan Review to discuss outdoor air calcs. They have reported that we did not use the correct tables for our ventilation tab. Actually I doubled check this and it is the EXACT table he listed. Perhaps they are misconstruing our use of the transient occupancy rule in both the code and the ASHRAE guidelines. This is where I said the owner may have to provide a letter stating that the space is actually only used at full capacity for under 3 hours on any given day – it is OVERFLOW not full use. Bob Martin called me back yesterday and I am calling him back right now to resolve.

06/15/06 – Bids were opened. Mechanical and Electrical are at budget. Once GC bid and is acceptable also. Partners are waiting on a formal approved bid tab to write the contracts for this project

05/25/06 – Drawings arrived today for this project.

05/22/06 – Spencer has submitted the drawings for bidding. Spencer and TTC discussed bidding estimates and the cost were entered and were below the original estimate. QCR called to say that 2 of TEI dwgs were mislabeled as mechanical dwgs when they were really electrical.

04/11/06 – TTC & Sam discussed project design today in office for layout. Sam met with Mike Bartlett yesterday to find a suitable HHW source for reheat. HHW was so far away that it doesn’t make sense for the space design we are using. TTC discussed zoning of the space today with Mike and decided the best and easiest way to serve these spaces based on our learning’s in the 450 Mediation space is to provide two zones with averaging thermostats to keep areas about the same temperature. Ease of use, easy to maintain and both zones can be kept in the MER adjacent to the air handler for ease of maintenance as well. TTC has to confirm with Mark Donnelly and Mike Bartlett economizers for this air handler on Wednesday.

04/11/06 – Final design is being put on the drawings. It appears that electric reheat will be used. TEI will break the room up into 2 zones total. One for mediation and one for the Jury commission. We will use averaging thermostats in the different rooms to adjust the air flow and temperature of the unit serving the space. This is a modular unit with electric reheat on the VAV boxes. A cooling coil and steam coil will provide the space and supply temperature to the main ducts. The two VAV boxes can be located in the new MER we are using for this job.

04/05/06 – TEI has been coordinating with CR and the owner extensively on the requirements of the new space. An owner provided sketch has created a higher density floor plan that Spencer sent to everyone. CR and TEI decided that ductwork can be ran in the space if the ceiling is lowered to 8′-0″. This will allow TEI to run supply and return ductwork in the interstitial 18″ of open space. A modular unit with direct outdoor air can be used to provide VAV boxes with supply air for the space. The radiators can be used for heating in the winter time. Mike pointed out that the reheat system is off during the summer so no reheat is available in the space. TEI will take this into consideration when designing the new system.

03/14/06 – Second Design meeting yielded a more formalized plan. CR will provide a mirrored floor plan (MS suggestion) and TEI will have mechanical and electric work done by April 14th for review with the owner. TEI will most likely replace the lights and the HVAC unit on this project.

03/07/06 – First Design Meeting. Scope defined for space but we need direction from owner on mediation’s part of the space and actual tenants requirements.

02/16/06 – SOSS sent but not approved.